OpenStudy (anonymous):

Using the Squeeze theorem, show that Limit x->0 x^2cos(1/x)=0

OpenStudy (anonymous):

Thankfully, the range of \(\cos x\) is bounded between \([-1,1]\), so we can assert that the value of \(x^2\cos\frac1x\) is between \(-x^2\) and \(x^2\). As \(x\) approaches \(0\), the values of the boundaries are both \(0\), and thus the only possible value for \(\lim_{x\to0}x^2\cos\frac1x\) is \(0\).

OpenStudy (anonymous):

u need to show with a graph or/and showing as piecewise funtion

OpenStudy (anonymous):

 Haha, shoddy drawing aside, the parabolic functions are the boundaries, whereas the actual function is constrained by said boundaries. At \(0\) there is only one possible region for the actual function to pass through,